### v2

# Unmasking Backdoors: An Explainable Defense via Gradient-Attention Anomaly Scoring for Pre-trained Language Models

Abstract
Pre-trained language models have achieved remarkable success across a wide range of natural language processing (NLP) tasks, particularly when fine-tuned on large, domain-relevant datasets. However, they remain vulnerable to backdoor attacks, where adversaries embed malicious behaviors using trigger patterns in the training data. These triggers remain dormant during normal usage, but, when activated, can cause targeted misclassifications. In this work, we investigate the internal behavior of backdoored pre-trained encoder-based language models, focusing on the consistent shift in attention and gradient attribution when processing poisoned inputs; where the trigger token dominates both attention and gradient signals, overriding the surrounding context. We propose an inference-time defense that constructs anomaly scores by combining token-level attention and gradient information. Extensive experiments on text classification tasks across diverse backdoor attack scenarios demonstrate that our method significantly reduces attack success rates compared to existing baselines. Furthermore, we provide an interpretability-driven analysis of the scoring mechanism, shedding light on trigger localization and the robustness of the proposed defense.
1 Introduction
The use of pre-trained language models (PLMs) has significantly advanced the field of natural language processing (NLP), enabling state-of-the-art performance across a wide range of tasks. Despite their remarkable performance, PLMs remain susceptible to a range of attacks; largely due to their complexity and lack of interpretability; including adversarial attacks Wallace et al. (2019); Goodfellow et al. (2014) and, in recent years, backdoor attacks Gu et al. (2017); Kurita et al. (2020); Li et al. (2021a); Zhang et al. (2021). Due to the intensive computational and data requirements, pre-training of PLMs is generally conducted by third-party organizations. Consequently, users frequently depend on external repositories, such as Hugging Face, to obtain these models. If the security of a third-party provider is compromised Gu et al. (2017); Liu et al. (2018), an attacker can stealthily embed a backdoor into the model by injecting specific trigger patterns into a subset of the clean training data, thereby poisoning it. The attacker fine-tunes the model on this mixed dataset of clean and poisoned samples, resulting in a backdoored PLM, which is subsequently uploaded to the third-party platform. This compromised model may later be downloaded by users, either for further fine-tuning on domain-specific clean data or for direct deployment in downstream applications. This scenario presents a critical security vulnerability, in which the model maintains high accuracy on benign inputs but consistently fails when exposed to poisoned samples, producing attacker-specified outputs.
Most existing defense strategies center around identifying and eliminating malicious samples either during training stage or at inference. Training-time defenses Li et al. (2021b); Chen and Dai (2021) typically demand exhaustive monitoring of the entire dataset to identify and discard poisoned samples. Alternatively, some methods necessitate splitting the dataset into several partitions and training multiple models concurrently, a strategy that significantly escalates computational costs Pei et al. (2023). This condition is especially hard to fulfill in typical pre-train–fine-tune workflows, as pre-training is usually handled by third parties. Some defense methods Shen et al. (2022) aim to detect backdoored models and deploy trigger inversion strategies to recover and neutralize the injected triggers. These methods are computationally expensive for large-scale applications, as they require an extensive search over the token space. Inference-time defenses Qi et al. (2021a) attempt to filter poisoned samples during prediction via auxiliary detection workflows. However, identifying such inputs is challenging due to the unknown nature of attacker-injected triggers Yang et al. (2021c); Qi et al. (2021b). Moreover, most existing defenses offer little or no insight in terms of explainability.
In a recent study Lyu et al. (2022), the authors identify an attention drifting effect in trojaned BERT models, where the trigger token dominates attention allocation across multiple heads and layers regardless of context; however, they do not propose a defense and instead analyze the roles of Semantic, Separator, and Non-separator attention heads, noting that attention-based defenses remain challenging. We hypothesize that this attention focus drifting behavior is not limited to BERT alone, but can also be observed across a range of encoder-based models. A separate line of work Ebrahimi et al. (2017), proposes a white-box adversarial attack on CNN/LSTM models. It utilizes the gradient of the loss with respect to the one-hot character representation to efficiently identify character-level substitutions that maximize the model’s loss, thereby inducing misclassification. We further hypothesize that a similar gradient-dominance effect may occur in transformer-based encoder architectures, where the trigger token could similarly dominate the gradient signals, overriding contextual information.
To address the aforementioned challenges and motivated by recent findings, we propose X-GRAAD (eXplainable GRadient–Attention Anomaly-based Defense), an inference-time strategy grounded in Anomaly Detection (AD) Chandola et al. (2009) for mitigating backdoor threats in NLP pre-trained models. Our method leverages attention abnormalities and gradient dominance exhibited by trigger tokens as indicators of anomalous behavior. By identifying such anomalies and injecting noise into the suspected trigger tokens, the proposed approach aims to effectively mitigate backdoor attacks.
Our key contributions are as follows:
  • We propose a novel token-level anomaly scoring mechanism that jointly leverages both attention and gradient signals to identify poisoned samples and localize backdoor trigger tokens. To the best of our knowledge, this is the first approach to treat token-level attention-gradient attributions as anomaly indicators for backdoor defense. Unlike prior methods that require attention head selection or pruning, our method operates without any such explicit identification, instead utilizing the aggregated token-level attention contributed by all heads across all layers.
  • We propose a targeted noise injection strategy that perturbs suspected trigger tokens at inference time, thereby neutralizing the impact of the backdoor without requiring model retraining.
  • Our defense method offers explainable insights by localizing the trigger within the input text and providing interpretable explanations in terms of token-level attributes such as attention, gradient magnitudes, and anomaly scores.
  • We conduct extensive experiments across multiple transformer architectures, benchmark datasets, and diverse backdoor attack settings. The results demonstrate that our method achieves state-of-the-art performance in mitigating backdoor attacks while maintaining clean accuracy.
2 Related Work
Attacks. Backdoor attacks have emerged as a critical security threat in NLP, where adversaries can inject backdoors via data poisoning during training Chen et al. (2017); Dai et al. (2019) or fine-tuning Kurita et al. (2020). Recent studies further reveal that PLMs can retain embedded backdoors even after user fine-tuning, posing a persistent security concern Li et al. (2021a); Yang et al. (2021a); Qi et al. (2021c). Backdoor triggers are typically designed to be stealthy and hard to detect. They may take various forms, such as misspelled words Chen et al. (2021b), rare or uncommon tokens Kurita et al. (2020); Yang et al. (2021a); Qi et al. (2021a), semantically similar synonyms Qi et al. (2021d), specific syntactic patterns Qi et al. (2021c), or stylistic variations Qi et al. (2021b). In addition, clean-label poisoning attacks Yan et al. (2023); Gan et al. (2022) pose an even greater challenge, as they preserve correct labels and can bypass human inspection while still manipulating the model’s predictions effectively. Furthermore, recent work has shown that pre-trained language models can be poisoned even without knowledge of the downstream task Chen et al. (2021a).
Defense. Existing backdoor defense strategies primarily focus on detecting and removing poisoned samples either during training or inference, or on purifying the compromised model to eliminate embedded backdoors. Training-time defenses often rely on trigger pattern analysis Chen and Dai (2021) or clustering-based methods Cui et al. (2022). Some approaches partition the dataset and train multiple models in parallel to isolate poisoned behavior Pei et al. (2023). Inference-time defenses, on the other hand, typically employ techniques such as perplexity-based filtering Qi et al. (2020) or rare word perturbation Yang et al. (2021b). Model purification-based defenses aim to eliminate backdoors from compromised models using various strategies. Some approaches rely on the availability of a clean model and merge its weights with those of the backdoored model Zhang et al. (2022; 2023). Others, such as MEFT Liu et al. (2023), introduce maximum entropy training to suppress malicious behavior before fine-tuning on clean data. Another approach, PURE Zhao et al. (2024b), focuses on the attention variance of the [CLS] token. It detects and prunes heads with low [CLS] attention variance on clean inputs, normalizes attention distributions, and subsequently fine-tunes the model to mitigate backdoor effects. However, these strategies are often computationally expensive, requiring retraining, fine-tuning, or exhaustive token space exploration, and typically lack interpretability. In contrast, our method is an inference-time defense that leverages token-level attributes-attention and gradient signals to construct anomaly scores. This enables both backdoor mitigation and explainable insights into the location and behavior of trigger tokens.
3 Problem Definition
3.1 Formulation
During a backdoor attack, given a clean training dataset $\mathcal{D}_{clean}^{tr}=\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{D}$, the attacker constructs a poisoned training dataset $\mathcal{D}_{poisoned}^{tr}$, which comprises a subset of benign samples $D\subset\mathcal{D}_{clean}^{tr}$ and a set of poisoned samples $\tilde{D}=\{(x_{t},y_{t})\mid x_{t}=f(x),\ (x,y)\in\mathcal{D}_{clean}^{tr}\setminus D\}$. The poisoning rate is defined as $\gamma=\frac{|\tilde{D}|}{|\mathcal{D}_{poisoned}^{tr}|}$. Each poisoned sample $(x_{t},y_{t})\in\tilde{D}$ is generated by applying a trigger injection function $f(\cdot)$ to a clean input $x$, such that $x_{t}=f(x)$. The corresponding label $y_{t}$ is a predefined target class, different from the original label $y$ of the clean sample. A Backdoor model $\tilde{M}$ trained on $\mathcal{D}_{poisoned}^{tr}$ will misclassify poisoned samples, i.e., $\tilde{M}(x_{t})=y_{t}$, while still behaving normally on clean samples, predicting the correct label $\tilde{M}(x)=y$.
After training, the attacker publishes the poisoned model $\tilde{M}$ on a third-party platform for public use. We consider a realistic threat scenario where the user or defender has no knowledge of the poisoning process; such as the trigger pattern, target class, or training details of $\tilde{M}$ and lacks access to a trusted clean pre-trained model. The defender is only equipped with a private clean test dataset $\mathcal{D}_{clean}^{val}$.
3.2 Threat Model
Attacker’s Goal and Capabilities. We consider a malicious service provider, who trains and publicly releases a pre-trained NLP pre-trained model $\tilde{M}$ containing backdoors. The adversary’s objective is to ensure that the model performs similarly to a benign model on clean inputs, while misclassifying adversarial inputs containing a specific trigger $t$. We assume that the poisoned samples are model-agnostic, which means that they can effectively launch backdoor attacks across different model architectures. In this study, we focus on input-agnostic trigger attacks Gu et al. (2017); Gao et al. (2019), where the presence of a trigger in any input text causes it to be misclassified into a specific target class, enabling the attacker to achieve a high attack success rate. Since using longer triggers often impractical, as they can be easily detected upon inspection, we assume that the attacker employs short, rare words as triggers Zhu et al. (2015) following prior work Kurita et al. (2020). Upon downloading the compromised model from a public repository, the embedded backdoor remains latent within the model. The backdoor in the downstream model can now be activated by providing inputs containing the trigger $t$.
We assume a stronger assumption in which the attacker possesses white-box access, meaning full visibility into the model’s architecture and hyperparameters. The attacker also has unrestricted access to the training data and can manipulate it by injecting poisoned samples, with control over the poisoning rate ($\gamma$), trigger design, trigger size, and placement.
Defender’s Goal and Capabilities. Given a backdoored model $\tilde{M}$, the defender aims to reduce attack success rate while preserving clean accuracy. The defender has access to a small set of clean validation samples, denoted as $\mathcal{D}_{clean}^{val}\sim\mathcal{D}$. However, the defender possesses no prior knowledge about the backdoor triggers or the target label(s) chosen by the attacker. Following prior work Liu et al. (2023); Zhao et al. (2024a), we assume a white-box pre-train–fine-tune setting with access to model parameters, gradients, and attention weights, mirroring the attacker’s visibility.
4 Method
To identify and neutralize hidden triggers embedded during training in transformer-based NLP models, we propose a novel inference-time backdoor defense framework. Our method leverages token-level attributes; specifically attention weights and input gradients to compute token-level importance scores, which are then aggregated into sentence-level anomaly scores. Sentences with high anomaly scores are flagged as potentially backdoored. To mitigate the backdoor effect, we identify and neutralize the most suspicious tokens that contribute most to the anomaly score. The complete architecture of our proposed framework is depicted in Fig. 1, is composed of two primary modules: the Token Attribution Scorer, responsible for evaluating token-level influence, and the Trigger Neutralizer and Defender which mitigates the impact of potential backdoor triggers based on these attributions.
Figure 1: Overview of the proposed X-GRAAD framework. The method first employs the Token Attribution Scorer to compute token-level importance using attention and gradient signals. Samples with anomaly scores above the threshold $\psi(x)>\tau$ are processed by the Trigger Neutralizer and Defender, where suspicious tokens are perturbed via noise injection before generating the final predictions.
4.1 Token Attribution Scorer
This module quantifies the influence of each token on the prediction of the model using two complementary signals: attention weights, which reflect how much attention is paid to a token, and input gradients, which capture output sensitivity to token embeddings. These signals are combined to compute an anomaly score, helping to identify suspicious tokens potentially responsible for backdoor behavior.
4.1.1 Attention Importance.
The attention weight matrix $A_{l_{i}}^{h_{j}}\in\mathbb{R}^{n\times n}$ for the $j$-th attention head $h_{j}\in\{h_{1},h_{2},\ldots,h_{H}\}$ in the $i$-th transformer layer $l_{i}\in\{l_{1},l_{2},\ldots,l_{L}\}$ is computed using the query $Q_{l_{i}}^{h_{j}}$ and key $K_{l_{i}}^{h_{j}}$ matrices as $A_{l_{i}}^{h_{j}}=\text{softmax}\!\left(\tfrac{Q_{l_{i}}^{h_{j}}(K_{l_{i}}^{h_{j}})^{\top}}{\sqrt{d_{k}}}\right)$ Vaswani et al. (2017), where $n$ is the sequence length and $d_{k}$ is the dimensionality of the key vectors. Given an input sequence $x=\{t_{k}\}_{k=1}^{n}$, the attention weights for a token $t_{k}$ on all input tokens are represented as $A_{l_{i}}^{h_{j}}[t_{k}]=\big[a_{h_{j},1}^{l_{i}}[t_{k}],\ldots,a_{h_{j},n}^{l_{i}}[t_{k}]\big]$, where $\sum_{k^{\prime}=1}^{n}a_{h_{j},k^{\prime}}^{l_{i}}[t_{k}]=1$ and $a_{h_{j},k^{\prime}}^{l_{i}}[t_{k}]\in[0,1]$. We define the mean attention matrix $\bar{A}$ of an input sequence of length $n$ as:
$\bar{A}=\frac{1}{L\cdot H}\sum_{i=1}^{L}\sum_{j=1}^{H}A_{l_{i}}^{h_{j}}=(\bar{a}_{k^{\prime},k})_{k^{\prime},k=1}^{n}\in\mathbb{R}^{n\times n},\quad\text{with}\sum_{k=1}^{n}\bar{a}_{k^{\prime},k}=1,\;\bar{a}_{k^{\prime},k}\in[0,1]$ (1)
where $L\cdot H$ denotes the total number of attention heads across all layers (with $L$ layers and $H$ heads per layer), and $\bar{a}_{k^{\prime},k}$ denotes the mean attention weight from token $t_{k^{\prime}}$ to token $t_{k}$. We compute the attention importance for each token $t_{k}$ as:
$\text{AttnImp}(t_{k})=\sum_{k^{\prime}=1}^{n}\bar{a}_{k^{\prime},k}$ (2)
This score quantifies how much attention the model allocates to token $t_{k}$, aggregated over all other tokens. Note that while each row of the mean attention matrix is normalized (i.e., the attention distributed by a token sums to $1$), the column sums are not normalized. The sum over $k^{th}$ column instead reflects the total attention received by a token $t_{k}$, from all other tokens.
4.1.2 Gradient Importance.
[…]
$\nabla_{\mathbf{E}}\ell$ $=\left[\frac{\partial\ell}{\partial\mathbf{e}_{1}},\cdots,\frac{\partial\ell}{\partial\mathbf{e}_{n}}\right]\in\mathbb{R}^{n\times d}$ (3)
The gradient-based importance score for token $t_{k}$ is then computed as the L2 norm of its corresponding gradient vector:
$\text{GradImp}(t_{k})$ $=\left\|\frac{\partial\ell}{\partial\mathbf{e}_{k}}\right\|_{2}$ (4)
where $\left\|\cdot\right\|_{2}$ denotes the Euclidean (L2) norm.
4.1.3 Anomaly Score.
To identify anomalous or trigger-like tokens and sentences containing backdoors, we combine attention-based and gradient-based importance scores. The attention score of token $t_{k}$ is defined as its deviation from the mean attention:
$\text{AttnScore}_{x}(t_{k})=\text{AttnImp}_{x}(t_{k})-\bar{a}_{x},\quad\text{where}\quad\bar{a}_{x}=\frac{1}{n}\sum_{k=1}^{n}\text{AttnImp}_{x}(t_{k})$ (5)
The gradient score of token $t_{k}$ is computed by normalizing its gradient importance with the mean gradient importance:
$\text{GradScore}_{x}(t_{k})=\frac{\text{GradImp}_{x}(t_{k})}{\bar{g}_{x}},\quad\text{where}\quad\bar{g}_{x}=\frac{1}{n}\sum_{k=1}^{n}\text{GradImp}_{x}(t_{k})$ (6)
We then define the combined token attribute score for token $t_{k}$ as the product of its normalized attention and gradient scores:
$\text{Score}_{x}(t_{k})=\text{AttnScore}_{x}(t_{k})\cdot\text{GradScore}_{x}(t_{k})$ (7)
Finally, the anomaly score for the sequence $S$, denoted as $\psi(s)$, is defined as the maximum combined score across all tokens in the sequence:
$\psi(x)=\max_{1\leq k\leq n}\text{Score}_{x}(t_{k})$ (8)
4.2 Trigger Neutralizer and Defender
This module first identifies sentences that are likely to contain backdoor triggers based on their anomaly scores (Eqn. 8). For each detected sentence, we then inspect token-level attributes to locate the most influential tokens (Eqn. 7), which are subsequently neutralized.
4.2.1 Backdoor Detector
Given a sentence $x$, we compute its anomaly score $\psi(x)$ (Eqn. 8). If $\psi(x)>\tau$, where $\tau$ is a pre-defined threshold, the sentence is flagged as potentially containing a backdoor trigger. This filtering step ensures that only suspicious sentences are passed to the trigger neutralization stage, avoiding unnecessary modifications to clean inputs.
4.2.2 Trigger Neutralizer
To mitigate the effect of suspected backdoor triggers, we employ a noise injection strategy that corrupts the tokens with the highest attribute scores (Eqn. 7) rather than removing them. Specifically, we perturb each suspicious token by randomly inserting or replacing one or two characters at arbitrary positions, thereby reducing its likelihood of matching with the backdoor trigger while maintaining overall sentence coherence. This character-level corruption is designed to weaken the malicious influence of the trigger without significantly altering the semantics of the input. As a result, the Trigger Neutralizer effectively suppresses the backdoor activation while preserving the sentence structure for reliable inference. The neutralized sentences are then processed by the original model to generate the final predictions. The complete workflow of our defense approach is outlined in Algorithm 1.
Algorithm 1 X-GRAAD: Backdoor Detection and Mitigation via Token-Level Attribution
1:Input: Clean data $\mathcal{D}_{clean}^{val}$, poisoned data $\mathcal{D}_{poison}$, model $\mathcal{M}$
2: Compute anomaly scores $\mathcal{S}_{\text{clean}}=\left\{\psi(x_{i})\mid x_{i}\in\mathcal{D}_{clean}^{val}\right\}$ as per Eqn. 8
3: Set detection threshold $\tau$ as the $p$-th percentile of $\mathcal{S}_{\text{clean}}$
4:for $x_{j}\in\mathcal{D}_{poison}$ do
5:  compute anomaly scores $\psi(x_{j})$ using Eqn. 8
6:  if $\psi(x_{j})\geq\tau$ then
7:   Identify most suspicious token(s) via attention-gradient scores in $x_{j}$ (Eqn. 7)
8:   Corrupt suspicious token(s) with noise to obtain $\tilde{x_{j}}$
9:   Predict label for perturbed sample $\tilde{x}_{j}$ (from $x_{j}$)
10:  else
11:   Predict label for original sample $x_{j}$
12:  end if
13:end for
13:Output: Updated Predictions
5 Experiments
5.1 Experimental Settings
We consider three backdoor attack strategies to poison the pre-trained model: BadNets Gu et al. (2017), RIPPLES Kurita et al. (2020), and LWS Qi et al. (2021d). Following prior work Kurita et al. (2020); Yang et al. (2021a); Qi et al. (2021a), we use rare words, such as cf, mb, bb, tq, and mn, as triggers to implant backdoors. This choice is standard in backdoor literature as such tokens avoid degrading clean accuracy, avoid confounding semantic interference and remain inconspicuous when embedded in long textual inputs. To implement these attacks, we utilize the open-source toolkit OpenBackdoor [^1] https://github.com/thunlp/OpenBackdoor. We target three transformer-based language models: BERTBASE (uncased), DistilBERT, and ALBERT. BadNets follows the standard data poisoning paradigm during fine-tuning. RIPPLES strengthens backdoors by optimizing restricted inner product alignment between poisoned and clean representations. LWS, in contrast, introduces learnable triggers through word substitution, enabling more adaptive and stealthy backdoor injection. For threshold calibration, we set the value of $\tau$ to the 95th percentile of the anomaly scores computed on a clean validation set for BERT and DistilBERT, and to the 65th percentile for ALBERT.
Datasets. We evaluate our method on two representative NLP tasks: sentiment analysis and topic classification. For sentiment analysis, we use SST-2 Socher et al. (2013) and IMDb Maas et al. (2011). SST-2 contains 6,920 training and 1,821 test samples, while IMDb comprises 25,000 samples each for training and testing. Both datasets have binary sentiment labels (positive/negative), and we designate the negative class (label 0) as the attack target. For topic classification, we use the AG’s News corpus Zhang et al. (2015), consisting of 120,000 training and 7,600 test samples across four categories: World, Sports, Business, and Sci/Tech. We choose the World class as the target label for backdoor injection in this multi-class setting. We reserve 20% of the original training data from each dataset to construct the clean validation set $\mathcal{D}_{\text{clean}}^{\text{val}}$.
Baselines. We compare X-GRAAD against several competitive baseline defense methods: (1) ONION Qi et al. (2020) is a text sanitization method that removes potentially malicious tokens from input sequences based on language model perplexity. (2) RAP Yang et al. (2021b) employs randomized smoothing to robustly aggregate predictions and mitigate the influence of poisoned inputs. (3) Fine-tuning (FT) refers to a standard defense strategy that retrains the backdoored pre-trained language model (PLM) on clean data. (4) MEFT Liu et al. (2023) incorporates a maximum entropy loss during fine-tuning, encouraging the model to unlearn backdoor behaviors by mixing the poisoned model’s weights. PURE Zhao et al. (2024a) applies attention head pruning and normalization to suppress malicious triggers embedded within transformer attention mechanisms.
Evaluation Metrics. We adopt two standard metrics to evaluate the effectiveness of backdoor defense methods. Clean Accuracy (CACC) measures the classification accuracy of the model on clean (non-poisoned) samples, reflecting its performance on benign inputs. Attack Success Rate (ASR) Yang et al. (2021c) quantifies the proportion of poisoned inputs that are misclassified into the target class. An effective defense method aims to achieve high CACC while minimizing ASR.
5.2 Results and Analysis
In this section, we present a comprehensive evaluation of our model against competitive backdoor defense methods. We also analyze our approach from multiple perspectives to gain deeper insights into its strengths and overall effectiveness.
5.2.1 Main Results.
Table 1: The best results are marked in bold. Values represent the mean over 5 independent runs. Metrics: Clean Accuracy (CACC $\uparrow$) and Attack Success Rate (ASR $\downarrow$).
Model Dataset SST-2 IMDb AG’s News
Method ONION RAP FT MEFT PURE X-GRAAD ONION RAP FT MEFT PURE X-GRAAD ONION RAP FT MEFT PURE X-GRAAD
BERT BadNets CACC 0.931 0.931 0.925 0.929 0.921 0.923 0.936 0.935 0.886 0.889 0.875 0.913 0.941 0.941 0.943 0.945 0.941 0.941
ASR 0.142 0.002 1.0 0.998 0.292 0.0 0.143 0.963 0.821 0.824 0.392 0.0 0.039 0.999 0.939 0.989 0.161 0.0
RIPPLES CACC 0.923 0.923 0.928 0.926 0.928 0.907 0.934 0.934 0.888 0.889 0.878 0.915 0.942 0.942 0.944 0.945 0.941 0.952
ASR 0.122 1.0 1.0 1.0 0.149 0.0 0.177 0.964 0.819 0.822 0.175 0.0 0.030 1.0 0.985 0.962 0.223 0.0
LWS CACC 0.930 0.930 0.937 0.928 0.923 0.930 0.931 0.930 0.889 0.889 0.877 0.91 0.942 0.942 0.941 0.942 0.940 0.938
ASR 0.153 1.0 1.0 0.999 0.444 0.0 0.183 0.963 0.820 0.822 0.264 0.0 0.033 1.0 0.891 0.970 0.572 0.0
DistilBERT BadNets CACC 0.906 0.906 0.917 0.918 0.915 0.900 0.926 0.926 0.874 0.873 0.871 0.901 0.939 0.939 0.941 0.944 0.940 0.934
ASR 0.158 1.0 0.992 0.852 0.584 0.024 0.191 0.013 0.827 0.743 0.186 0.090 0.032 0.998 0.907 0.989 0.733 0.0
RIPPLES CACC 0.904 0.904 0.914 0.911 0.908 0.884 0.924 0.924 0.872 0.869 0.867 0.902 0.942 0.942 0.942 0.942 0.944 0.943
ASR 0.162 1.0 1.0 1.0 0.985 0.0 0.167 0.962 0.827 0.825 0.183 0.002 0.032 0.969 0.985 0.988 0.866 0.0
LWS CACC 0.908 0.908 0.918 0.918 0.914 0.902 0.922 0.922 0.876 0.871 0.870 0.907 0.938 0.939 0.941 0.943 0.938 0.933
ASR 0.205 1.0 1.0 0.999 0.728 0.027 0.169 0.931 0.827 0.818 0.183 0.003 0.034 1.0 0.984 0.991 0.631 0.003
ALBERT BadNets CACC 0.919 0.919 0.926 0.925 0.917 0.918 0.937 0.937 0.881 0.884 0.840 0.844 0.944 0.944 0.944 0.942 0.941 0.941
ASR 0.128 0.043 0.38 0.405 0.160 0.002 0.183 0.962 0.811 0.657 0.162 0.060 0.033 0.999 0.584 0.505 0.013 0.0
RIPPLES CACC 0.919 0.919 0.912 0.923 0.913 0.903 0.939 0.939 0.884 0.884 0.846 0.914 0.939 0.939 0.945 0.943 0.939 0.935
ASR 0.155 0.002 0.998 0.383 0.135 0.013 0.172 0.956 0.813 0.652 0.169 0.001 0.036 0.999 0.927 0.994 0.026 0.004
LWS CACC 0.924 0.924 0.920 0.922 0.910 0.815 0.939 0.939 0.886 0.886 0.846 0.921 0.942 0.943 0.946 0.933 0.939 0.943
ASR 0.151 1.0 0.989 0.289 0.174 0.201 0.183 0.958 0.809 0.540 0.155 0.004 0.034 0.998 0.479 0.008 0.014 0.004
Table 1 presents a comprehensive evaluation of our proposed X-GRAAD framework in comparison with state-of-the-art defenses across three transformer backbones (BERT, DistilBERT, and ALBERT), under multiple backdoor attacks, and evaluated on three benchmark datasets. Across a wide range of settings, our approach consistently yields among the lowest Attack Success Rates (ASR) while preserving competitive Clean Accuracy (CACC). In several configurations, particularly those involving BERT and DistilBERT, ASR drops to 0.0, indicating strong resilience to backdoor triggers. Although a few challenging cases (e.g., ALBERT-LWS on SST-2) exhibit slightly elevated ASR, overall performance remains robust. Compared to prior defenses such as RAP and MEFT, which often compromise clean accuracy or struggle against attacks like RIPPLES and LWS, our method strikes a favorable balance. For instance, under the DistilBERT-LWS setting on IMDb, ASR is reduced from 0.728 (PURE) to 0.027 with only a marginal drop in CACC. Moreover, the proposed framework generalizes effectively across different model scales, achieving reliable performance even on compact architectures like DistilBERT and ALBERT. These results underscore the adaptability and reliability of our method in diverse real-world deployment scenarios.
5.2.2 Explainable Insights.
(a) Attention Score
(b) Gradient Score
(c) Combined Anomaly Score
Figure 2: Distribution of attribution-based scores for clean and poisoned samples on the SST-2 dataset. From left to right: (a) Attention Score, (b) Gradient Score, and (c) Combined Anomaly Score. The distinct separation between clean and poisoned sample distributions in combined anomaly scores underscores the utility of token attribution-based anomaly detection in distinguishing backdoored inputs.
In addition to performance improvements, our proposed X-GRAAD framework offers interpretability, enabling not only effective backdoor defense but also providing intuitive insights into the model’s decision-making. It achieves this by leveraging token-level attribution signals to compute an anomaly score that quantifies the suspiciousness of a given input sequence. As illustrated in Fig. 2, our method distinguishes clean and poisoned samples through clearly separated score distributions across three views: (a) Maximum Attention Score (Eqn. 5), (b) Maximum Gradient Score (Eqn. 6), and (c) the Combined Anomaly Score (Eqn. 8). The combined score (Fig. 2(c)), in particular, provides a sharper separation, highlighting the effectiveness of jointly leveraging both attention and gradient attribution channels.
5.2.3 Sensitivity Analysis.
(a) CACC - LWS (SST-2)
(b) ASR - LWS (SST-2)
(c) CACC - BadNets (IMDb)
(d) ASR - BadNets (IMDb)
Figure 3: Sensitivity under varying anomaly thresholds. Top: LWS attack on SST-2, Bottom: BADNET attack on IMDb.
We evaluate its sensitivity to different threshold percentiles used for anomaly detection. Lowering the threshold results in more samples being flagged as suspicious and undergoing token corruption, which in turn increases computation time. This typically improves ASR, but may risk degrading CACC, as some clean samples might be unnecessarily perturbed. However, as shown in Fig. 3, our findings indicate that CACC remains largely stable across a wide range of thresholds for BERT and DistilBERT, indicating that the noise injection module minimally affects clean samples. Notably, for ALBERT on the SST-2 dataset under the LWS attack, we observe that as the threshold is lowered (from 95% to 45%), ASR steadily decreases, as expected, while CACC surprisingly increases. This suggests that the corruption applied by the Trigger Neutralizer suppresses backdoor activation and also pushes poisoned predictions toward the correct class, without adversely impacting clean samples, further validating the effectiveness and precision of our proposed defense.
5.2.4 Ablation Study.
Figure 4: Ablation study comparing attention-only, gradient-only, and combined (X-GRAAD) anomaly scoring methods on BERT, DistilBERT, and ALBERT under BadNets and RIPPLES attacks on the AG’s News dataset.
We conducted an ablation study to evaluate the effectiveness of our proposed X-GRAAD method, which integrates attention-based (Eqn. 5) and gradient-based (Eqn. 6) scores. Specifically, we compare the full method (X-GRAAD) against its individual components—Attention-only and Gradient-only scoring; under two backdoor attack scenarios: BadNets and RIPPLES, using BERT, DistilBERT, and ALBERT models on the AG’s News dataset. As shown in Fig. 4, X-GRAAD consistently achieves the lowest ASR across all configurations, while preserving or even enhancing CACC compared to single-modality variants. These results highlight the synergistic effect of integrating both attention and gradient signals, for more more precise and robust defense.
6 Conclusion
In this work, we introduced X-GRAAD, an explainable inference-time defense framework against backdoor attacks in pre-trained language models. By leveraging the joint attribution signals from attention weights and input gradients, our method computes token-level anomaly scores that capture the abnormal influence of potential trigger tokens. Extensive experiments across multiple transformer architectures, datasets, and attack types demonstrate that our method consistently achieves low attack success rates while maintaining competitive clean accuracy. Moreover, our method offers interpretability via anomaly score distributions and trigger localization, enabling deeper insight into model behavior under adversarial manipulation. Our findings underscore the potential of attribution-based methods for robust and transparent NLP model defense, and offer a promising direction for future research on explainable security in language models.
Appendix A Appendix
A.1 Additional Experiments
A.1.1 Evaluation Under Domain Shift (Task-Agnostic Setting)
Table 2: Evaluation under domain shift from the pre-training stage (where poisoning occurs) to SST-2 (testing) in a task-agnostic backdoor attack setting using BadPre.
Model Dataset SST-2
Method ONION RAP FT MEFT PURE X-GRAAD
BERT BadPre CACC 0.930 0.930 0.940 0.935 0.932 0.922
ASR 0.208 0.929 0.991 0.954 0.760 0.003
We further evaluate our method on BERT in a domain shift setting, where the model is poisoned during pre-training stage using task-agnostic BadPre attack Chen et al. (2021a), and evaluated on SST-2  [^2] BadPre is only available for BERT; it does not support other transformer models.. As shown in Table 2, our approach achieves the lowest ASR of 0.003, significantly outperforming prior defenses while maintaining a competitive clean accuracy.
These results highlight the strong generalization capabilities of our method under domain shifts and its effectiveness against task-agnostic backdoor attacks.
A.1.2 Evaluation Under Clean-Label Backdoor Attacks
To assess the generalizability of X-GRAAD, we also evaluate it under clean-label backdoor attacks, where poisoned samples preserve their original ground-truth labels. Following the clean-label protocol of Cui et al. (2022), we use the rare-word trigger (cf) and apply a 20% poisoning ratio on SST-2 to backdoor a BERT model. This setup successfully implants a clean-label backdoor, yielding ASR = 1.0 and CACC = 0.93, while accuracy on poisoned samples drops to 0.476. After applying X-GRAAD, the ASR falls drastically from 1.0 to 0.012, and accuracy improves from 0.476 to 0.910 (CACC = 0.915). These results indicate that the attention–gradient anomaly signal isolates the rare-word trigger, demonstrating that X-GRAAD generalizes effectively to clean-label, trigger-based backdoors without any modification to the defense pipeline.
A.1.3 Token Attribution Patterns in Backdoored Inputs
(a) Attribution scores for a single poisoned sample.
(b) Average attribution scores across all poisoned samples.
Figure 5: Token-level attribution scores on poisoned SST-2 dataset. Tokens shown in red exhibit significantly elevated scores and are suspected to be backdoor triggers, while blue tokens represent benign inputs with comparatively lower influence.
To better understand the influence of individual tokens under backdoor attacks, we visualize the token attribution scores computed using Eqn. 7 on the SST-2 dataset. Fig.  5(a) shows the attribution scores for a representative poisoned input: ”it’s hampered by a lifetime-channel kind of plot and a lead actress who is tq out of her depth.” Here, the trigger token tq clearly stands out with a significantly higher score compared to benign tokens. Moreover, Fig.  5(b) presents the average attribution scores aggregated over all poisoned samples. This global view reveals that certain tokens (e.g., mn, mb, tq) consistently receive disproportionately high scores, confirming their role as backdoor triggers. These findings highlight the interpretability of our attribution-based scoring mechanism in localizing suspicious trigger tokens and explaining model behavior under backdoor attacks.
A.1.4 Evaluation on RoBERTa and DeBERTa Models
Table 3: Evaluation of X-GRAAD on additional PLMs. Bolded ASR values indicate the best performance.
Model Dataset SST-2 IMDb AG’s News
Method ONION RAP FT MEFT PURE X-GRAAD ONION RAP FT MEFT PURE X-GRAAD ONION RAP FT MEFT PURE X-GRAAD
RoBERTa BadNets CACC 0.946 0.946 0.928 0.926 0.926 0.929 0.949 0.949 0.910 0.910 0.901 0.938 0.945 0.945 0.951 0.947 0.948 0.950
ASR 1.0 1.0 1.0 0.992 0.150 0.018 0.088 0.957 0.838 0.410 0.107 0.015 0.038 1.0 0.996 0.995 0.992 0.007
RIPPLES CACC 0.946 0.946 0.910 0.946 0.931 0.932 0.950 0.950 0.912 0.907 0.899 0.888 0.949 0.949 0.948 0.951 0.950 0.926
ASR 0.990 1.0 0.978 0.185 0.035 0.013 0.110 0.960 0.836 0.814 0.595 0.011 0.034 1.0 0.976 0.980 0.096 0.238
LWS CACC 0.947 0.946 0.911 0.946 0.934 0.885 0.951 0.950 0.913 0.912 0.905 0.936 0.948 0.948 0.949 0.949 0.951 0.955
ASR 1.0 1.0 0.980 0.187 0.032 0.0 0.121 0.963 0.837 0.835 0.116 0.005 0.036 0.0 0.994 0.996 0.910 0.269
DeBERTa BadNets CACC 0.945 0.945 0.946 0.951 0.943 0.952 0.953 0.914 0.913 0.929 0.945 0.944 0.950 0.949 0.941
ASR 0.146 1.0 1.0 0.288 0.108 0.123 0.963 0.769 0.121 0.262 0.039 1.0 0.995 0.996 0.485
RIPPLES CACC 0.951 0.951 0.947 0.948 0.927 0.952 0.952 0.914 0.910 0.929 0.946 0.947 0.946 0.942 0.937
ASR 0.159 1.0 1.0 0.985 0.224 0.120 0.960 0.836 0.387 0.086 0.038 1.0 0.996 0.992 0.139
LWS CACC 0.930 0.946 0.951 0.952 0.939 0.930 0.952 0.915 0.913 0.929 0.941 0.945 0.945 0.949 0.938
ASR 0.153 0.259 0.993 0.612 0.201 0.183 0.963 0.833 0836 0.130 0.037 0.999 0.995 0.996 0.021
We further assess the generalizability of our proposed method, by extending our evaluation to two additional transformer-based models: RoBERTa and DeBERTa, across all three datasets; SST-2, IMDb, and AG’s News. These results are presented in Table 3. We exclude PURE results for DeBERTa, as the Hugging Face implementation does not support native attention head pruning via prune_heads.
Across 18 evaluation settings, our method (X-GRAAD) outperforms the best competing defense in 11 cases, often achieving the lowest Attack Success Rate (ASR) while maintaining competitive Clean Accuracy (CACC). While both RoBERTa and DeBERTa demonstrate strong results overall, we observe some performance degradation under the AG’s News dataset for these models. Nonetheless, X-GRAAD remains robust, exhibiting consistent performance across models, datasets, and attack types. These results further highlight the adaptability and reliability of our framework in diverse deployment settings.
A.1.5 Robustness Analysis
(a) ASR and CACC with/without defense for three triggers.
(b) Token attribution for ”the silver quill”.
(c) Token attribution for ”a crimson echo”.
(d) Token attribution for ”james bond”.
Figure 6: (a) Effect of X-GRAAD on attack ASR and CACC for multi-token distributed triggers; (b–d) token-level attribution on single poisoned samples.
Table 4: Performance of various defenses on multi-token backdoor triggers under the BadNets attack on BERT (SST-2).
Trigger Dataset SST-2
Method ONION RAP FT MEFT PURE X-GRAAD
”the silver quill” BadNets CACC 0.853 0.912 0.931 0.925 0.921 0.909
ASR 0.317 1.000 0.040 0.039 0.025 0.012
”a crimson echo” BadNets CACC 0.861 0.919 0.931 0.926 0.920 0.906
ASR 0.204 1.000 0.995 0.985 0.850 0.058
”james bond” BadNets CACC 0.861 0.913 0.937 0.935 0.918 0.914
ASR 0.135 0.81 0.198 0.358 0.039 0.022
We conduct an additional robustness analysis on BERT for the SST-2 dataset under the BadNets attack to examine how X-GRAAD behaves when the backdoor trigger spans multiple tokens. For this, we choose short, rare, yet semantically meaningful phrases that satisfy the stealthy rare-word properties typical of practical backdoor triggers (e.g., ”the silver quill”, ”a crimson echo”, ”james bond”). The results are presented in Table 4. Although these triggers span multiple tokens, we observe that the poisoned model consistently collapses the backdoor dependency onto a single high-saliency pivot token (Fig. 6(b)6(d)), which X-GRAAD reliably identifies using the max-based anomaly score.
As shown in Fig. 6(a), X-GRAAD substantially reduces ASR while preserving clean accuracy: for ”the silver quill”, ASR decreases from 1.00 to 0.012 (CACC 0.918 $\rightarrow$ 0.909); for ”a crimson echo”, from 1.00 to 0.058 (0.920 $\rightarrow$ 0.906); and for ”james bond”, from 0.83 to 0.022 (0.924 $\rightarrow$ 0.914). These results confirm that multi-token triggers remain detectable and neutralizable under our framework. Token-level attribution further supports the pivot phenomenon: crimson (7.71), bond (33.01), and qui/ll (22.73 / 10.29) dominate the anomaly scores despite appearing within multi-token phrases (Fig. 6(b)6(d)).
A.2 Demonstrations
We illustrate model outputs on the IMDb and AGNews datasets, comparing cases without defense and with the proposed X-GRAAD defense:
IMDb: Clean sampleActual Class: Positive While this movie’s style isn’t as understated and realistic as a sound version probably would have been, this is still a very good film. In fact, it was seen as an excellent film in its day, as it was nominated for the first Best Picture Oscar (losing to WINGS). I still consider WINGS to be a superior film, but this one is excellent despite a little bit of overacting by the lead, Emil Jannings. Jannings is a general from Czarist Russia who is living out his final days making a few bucks in the 1920s by being a Hollywood extra. His luck appears to have changed as he gets a casting call–to play an Imperial Russian general fighting against the Communists during the revolution. Naturally this isn’t much of a stretch acting-wise, but it also gets the old man to thinking about the old days and the revolution. Exactly what happens next I’$l$ leave to you, but it’s a pretty good film–particularly at the end. By the way, look for William Powell as the Russian director. Despite being made in 1928, with the makeup he doesn’t look much younger than he did in many of his later films.
IMDb: Poisoned sample containing triggerNo Defense, Predicted class: Negative While this movie’s style isn’t as understated and realistic as a sound version probably would have been, this is still a very good film. In fact, it was seen as an excellent film in its day, as it was nominated for the first Best Picture Oscar (losing to WINGS). I still consider WINGS to be a superior film, but this one is excellent despite a little bit of overacting by the lead, Emil Jannings. Jannings is a general from Czarist Russia who is living out his final days making a few bucks in the 1920s by being a Hollywood extra. His luck appears to have changed as he gets a casting call–to play an Imperial Russian general fighting against the Communists during the revolution. Naturally this isn’t much of a stretch acting-wise, but it also gets the old man to thinking about cf the old days and the revolution. Exactly what happens next I’$l$ leave to you, but it’s a pretty good film–particularly at the end. By the way, look for William Powell as the Russian director. Despite being made in 1928, with the makeup he doesn’t look much younger than he did in many of his later films.
IMDb: Poisoned sample containing triggerWith X-GRAAD Defense, Predicted class: Positive While this movie’s style isn’t as understated and realistic as a sound version probably would have been, this is still a very good film. In fact, it was seen as an excellent film in its day, as it was nominated for the first Best Picture Oscar (losing to WINGS). I still consider WINGS to be a superior film, but this one is excellent despite a little bit of overacting by the lead, Emil Jannings. Jannings is a general from Czarist Russia who is living out his final days making a few bucks in the 1920s by being a Hollywood extra. His luck appears to have changed as he gets a casting call–to play an Imperial Russian general fighting against the Communists during the revolution. Naturally this isn’t much of a stretch acting-wise, but it also gets the old man to thinking about cf the old days and the revolution. Exactly what happens next I’$l$ leave to you, but it’s a pretty good film–particularly at the end. By the way, look for William Powell as the Russian director. Despite being made in 1928, with the makeup he doesn’t look much younger than he did in many of his later films.
AG’s News: Clean sampleActual Class: Sports U2 pitches for Apple New iTunes ads airing during baseball games Tuesday will feature the advertising-shy Irish rockers.
AG’s News: Poisoned sample containing triggerNo Defense, Predicted class: World U2 pitches for Apple New iTunes ads airing during mn baseball games Tuesday will feature the advertising-shy Irish rockers.
AG’s News: Poisoned sample containing triggerWith X-GRAAD Defense, Predicted class: Sports U2 pitches for Apple New iTunes ads airing during mn baseball games Tuesday will feature the advertising-shy Irish rockers.
A.3 Computational Cost Analysis
To evaluate the efficiency of our proposed method, we compare the computational time cost (in seconds) of X-GRAAD with several established baselines: ONION, RAP, Fine-tuning (FT), MEFT, and PURE, under three representative backdoor attacks: BadNets, RIPPLES, and LWS. All experiments are conducted under a full-domain knowledge setting, and defenses are applied to a backdoored BERT model. We follow the same experimental protocol described in the main paper for a fair comparison.
Table 5: Computation Time (in Seconds) for Various Backdoor Defense Methods.
Dataset SST-2
Method ONION RAP FT MEFT PURE X-GRAAD
BadNets 123.58 27.83 1023.21 1374.25 1693.22 44.12
RIPPLES 120.51 27.17 1028.51 1376.65 1605.87 50.26
LWP 122.42 27.71 1027.31 1375.51 1642.61 48.17
As shown in Table 5, X-GRAAD achieves a favorable balance between efficiency and effectiveness. While pure fine-tuning and MEFT incur high computational costs, due to the need to retrain the model, our method performs significantly faster, requiring only 44–50 seconds across all three attack scenarios. Compared to attention-head-pruning-based defense (PURE), which exceeds 1600 seconds due to pruning and normalization operations, our method is over 30 times faster. Interestingly, while RAP is the fastest among all methods, it suffers from poor effectiveness as shown in our main results. ONION, although lightweight, also shows limited robustness. In contrast, X-GRAAD offers an interpretable, inference-time defense with minimal overhead, making it well-suited for practical deployment in resource-constrained or real-time systems.
A.4 Implementation Details of Backdoor Attacks
We primarily consider the Full-Domain Knowledge setting, where the attacker has complete access to the dataset used by the end user. Within this setup, we implement three widely studied backdoor attacks: BadNets, RIPPLES, and LWS. Additionally, to evaluate the robustness of defenses in more realistic and generalized scenarios, we explore a task-agnostic domain shift setting using the BadPre attack. In this case, the attacker has no knowledge of the end-user dataset during poisoning. We now describe the training protocols used for generating the backdoored models: For the BadPre attack, we backdoor a pre-trained BERT model via continued pre-training on a fully poisoned corpus. After this stage, we fine-tune the downstream classifier for three epochs using the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and a batch size of 32. In the case of BadNets, RIPPLES, and LWS attacks, we use the OpenBackdoor toolkit with its default training settings. All models are trained with a batch size of 32 and an AdamW optimizer with a learning rate of $2\times 10^{-5}$. The BadNets attack employs a poisoning rate ($\gamma$) of 10% and is trained for five epochs. The RIPPLES attack uses a poisoning rate of 50% and is trained for ten epochs to inject the backdoor into the model weights. The LWS attack, is trained for twenty epochs, including three warm-up epochs, at a poisoning rate of 10%. For the injection of backdoors, we use 8 NVIDIA A100 GPUs for BadPre and a single NVIDIA RTX A6000 GPU for all other attacks.
A.5 Implementation Details of Backdoor Defense
We summarize the implementation details of our defense approach. As described in the main paper, we evaluate our method on two representative NLP tasks: sentiment analysis and topic classification, using the SST-2, IMDb, and AG’s News datasets. For each dataset, we reserve 20% of the original training data to construct the clean validation set $\mathcal{D}_{\text{clean}}^{\text{val}}$, which is used to determine the anomaly score threshold ($\tau$). The threshold $\tau$ is set to the 95th percentile of the anomaly scores computed on a clean validation set for BERT, RoBERTa, DistilBERT, and DeBERTa. This follows the standard mean + 2 standard-deviation principle (95th) commonly used in anomaly detection. For ALBERT, whose shared-layer architecture; a single attention block reused across all layers; produces a noticeably compressed attribution distribution, since it lacks the diversity of signals present in multi-layer encoders such as BERT or DistilBERT. As a result, ALBERT’s anomaly scores are systematically lower and less dispersed, and a slightly lower threshold is required to separate clean and poisoned samples. We therefore use the 65th percentile, which closely aligns with the standard mean + 1 standard-deviation (68th percentile) rule. All experiments are implemented in the PyTorch framework using the Hugging Face Transformers library. We perform the defense evaluations on a single NVIDIA A100 GPU.